OpenStudy (jylarai):

"Hue is arranging chairs. She can form 6 rows of a given length with 3 chairs left over, or 8 rows of that same length if she gets 11 more chairs. Write and solve an equation to find how many chairs are in that row length"

OpenStudy (anonymous):

let n be number of chairs in each row,then 6n+3=8n-11 8n-6n=3+11 2n=14 n=7 number of chairs=6n+3=?

OpenStudy (jylarai):

45 chairs?

OpenStudy (anonymous):

correct.